OpenStudy (anonymous):

The sum of an integer and its reciprocal is 65/8. Find the integer. If the problem has two solutions, type your answers in the box below with NO spaces and separating them with a comma. Example: 2,7

OpenStudy (fwizbang):

Call the number you are looking for x. The reciprocal of x is 1/x. When you add them together, you get x + 1/x = 65/8. Multiply both sides of this eqn by x to get x^2 +1 = 65/8 x move the 65/8 x to the left side and then solve for x using the quadratic formula.....

OpenStudy (anonymous):

8,1/8 ?

OpenStudy (fwizbang):

perfect, but 1/8 isn't an integer........

OpenStudy (anonymous):

so just 8?

OpenStudy (fwizbang):

yes